Yesterday in class I was told if you had a problem like figureA you put a one in front of the variable, in this case x. Today I had a test and I did just that. After the test he went over the answers, and now he says do it as in figureB. Which way is correct? I have another test tomorrow and don't want to make the same mistakes. Thanks for your time.
FigureA: 4+x=4+1x
FigureB: 4+x=4x
Sincerely,
Katie

Answer by prince_abubu(198)

Figure A: 4 + x = 4 + 1x.

That's so true. I guess your teacher is trying to point out that when you write a variable by itself, there really is a 1 being multiplied to it, but you don't really write it. The mere presence of a single variable automatically means that there's one of it. If anything else, you would have to say 2x if there were really 2 x's.

Figure B: 4 + x = 4x <----- INCORRECT

Four plus some number does not mean to multiply 4 to whatever that number is. That's like saying 5 + 6 = 5*6, which is obviously wrong.

I'm not sure what you mean by "do it as in figure B".